# Empowering Snapshot Compressive Imaging: Spatial-Spectral State Space Model with Across-Scanning and Local Enhancement

### v1

@@ 4.3 Real Data Results
Figure 5: Spectral Density Curves.
We further apply our method to reconstruct real HSIs. We retrain our model on the CAVE cave and KAIST kaist datasets, conducting tests on real scenes for experiments. Following the configurations established in MST mst , We introduce 11-bit shot noise during training to simulate real measurement conditions. As shown in Fig. 6, we compare the reconstructed images from two real scenes (Scene 3 and Scene 4) using our ASLE-SSM and nine SOTA methods. Our model achieved superior results when compared to other models. Specifically, In Scene 3, our image preserves the most information compared to the results of other models. A noticeable protrusion consistent with the original image can be observed in the center of our result and other methods fail to reproduce.
Additionally, our images do not exhibit significant artifacts or ghosting. In Scene 4, We can observe that the face shape in the original image is relatively rounded. However, the reconstruction results of previous methods generate an almond-shaped face similar to the compressed measurements, indicating that these methods did not learn this information. In contrast, our results clearly show a fuller face on both sides. Additionally, previous reconstruction methods often only reconstruct one side of the mouth, while our image noticeably reconstructs the whole mouth, though there is a slight misalignment between the two sides. Furthermore, our images do not exhibit nonexistent features that appear in previous methods’ results, such as noses, wrinkles, or gaps that are not present in the original RGB image. These results demonstrate that our ASLE-SSM surpasses other methods in reconstructing image details and suppressing noise in real scenes.
[Uncaptioned image] Real Measurement [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] RGB Ref. Meas CST-L DGSMP GAP-Net HDNet [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] HLRTF DiffSCI $\lambda$-Net MST-L TSA-Net ASLE-SSM-L
Refer to caption Real Measurement Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption RGB Ref. Meas CST-L DGSMP GAP-Net HDNet Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption HLRTF DiffSCI $\lambda$-Net MST-L TSA-Net ASLE-SSM-L
@@ 4.4 Ablation Study
Break-down Ablation. We conduct an ablation study by removing the LE-Spatial Mechanism and AS-Spectral-Spatial Mechanism from the ASLE-SSM network. The experiment results are shown in Table 3. Our goal is to explore the contributions of each mechanism to the results. The baseline barely adopts the U-shaped framework without any SSM modules.
Table 3: Ablation study of ASLE-SSM’s components
@@
Baseline w/o SSM 0.18 27.41 0.678
with LE-SSM 0.62 34.48 0.942
with AS-SSM 0.37 31.16 0.863
with LE + AS SSM 0.81 34.74 0.946
When we respectively employ LE-SSM and AS-SSM blocks, our model achieves 7.07 dB and 3.75 dB improvements in PSNR over the baseline, respectively. Employing both mechanisms simultaneously can further enhance the model’s performance. Here we need to point out that, in consideration of maintaining lightweight, we only utilized a single-direction scan in AS-SSM. In fact, we could also deploy a four-direction scanning similar to LE-SSM. However, each scan direction implies additional computational costs so we decided not to adopt this way.
Table 4: Comparisons of AS mechanism and common scanning.
@@
w/o Spectral SSM 34.48 0.942
with AS Mechanism 34.74 0.946
with Common Scanning 34.55 0.943
Across-Scanning Mechanism. To demonstrate the advantages of our across-scanning mechanism compared to common spectral scanning, we conduct the experiment on model performance under both methods. As illustrated in Table 4, when we adopt common scanning, it yields a 0.07 dB improvement. In contrast, when we apply the proposed across-scanning mechanism, our method achieves a significant increase of 0.26 dB compared to not using spectral SSM. Both methods were experimented with under the condition of a single scan. The superiority of our method mainly comes from paying more attention to adjacent bands and pixels, thus better extracting the local features of HSIs along both spatial and spectral dimensions. The results verify the role of our AS mechanism.
@@
w/o Spatial SSM 31.16 0.863
with LE Machenism 34.48 0.942
with Global Scanning 34.40 0.940
Local-Enhancement Mechanism. Additionally, we conduct an ablation study of our LE mechanism on the CAVE and KAIST datasets to assess its contributions. As shown in Table 5, The evaluations begin with the model without SSM in the spatial dimension, which served as our initial baseline with a PSNR of 31.16 dB. Specifically, the addition of the global four-direction scanning along the spatial context yields a significant increase of 3.32 dB in PSNR and 0.077 in SSIM, demonstrating its effectiveness in modeling the global spatial context.
We further apply our LE Machenism along the spatial dimension and achieve a 0.08 dB increase compared to the global scanning. our method only changes the scanning order and does not introduce any additional parameter consumption or computational cost to achieve a relatively better performance. In other words, we do not use CNN layers or spatial attention mechanisms as in other methods localmamba . In fact, Without increasing any additional cost and with the reconstructed results already achieving a relatively good level, 0.08 dB in PSNR and 0.002 in SSIM are significant improvements. The ablation experiment demonstrates the importance of local features in HSIs and verifies the effectiveness of our LE Mechanism.
@@ 5 Conclusion
In this paper, we address two limitations of previous SSM-based models by our proposed mechanisms: insufficient utilization of channel (spectral) information, and limited local detail extraction capability. By designing a novel across-scanning mechanism, we fully utilize the rich spectral information in HSIs to guide the image reconstruction process. At the same time, our scanning method based on patched data cubes pays more attention to the local similarity of HSIs between adjacent bands and pixels, achieving better results compared to simply scanning along the channel dimension. By introducing local enhancement in the Spatial SSM, we integrate spatial context from both global and local perspectives, achieving better results than using a global SSM alone.
Despite the significant progress of our method, we also acknowledge that as the first framework to adopt the SSM mechanism in HSI reconstruction, our network still has room for further improvement. For example, we could design a lightweight and effective way to fuse local and global branch information in the spatial dimension, rather than simply adding them together. However, the key challenge lies not in improving the effect itself, but in controlling computational costs. We will explore this aspect in the future with the aim of designing a more efficient network architecture.

### v2

@@ 4.3 Real Data Results
We apply our model and the proposed masked strategy to reconstruct real HSIs, retraining on the CAVE [38] and KAIST [10] datasets and simulating real conditions with 11-bit shot noise. Fig. 7 compares reconstructions from two real scenes using our original CS-Mamba model, the masked version, and four state-of-the-art methods. When enlarging the image results, we can clearly observe that the original CS-Mamba output exhibits a peculiar grid-like noise around the shoulders and face of the subject. In contrast, our mask method effectively mitigates this issue, producing a more visually pleasing effect compared to previous approaches, achieving smoother and more accurate reconstruction.
We further conducted a separate comparison between the original model and the model employing the masked strategy in Scene 1. As shown in Fig. 4, the original model exhibits some wrinkles and artifacts on the raised areas of the petals, with edges becoming blurred, and the stem displaying certain noise patterns. In contrast, the ground truth image presents smooth elevations. This issue is significantly alleviated with our masked method, obtaining a marked improvement in visual quality. Additional experiments will be included in the supplementary material to further substantiate the validity of our approach and conclusions, while Visual comparisons here sufficiently demonstrate the effectiveness of our masked training method.
Refer to caption Real Measurement Refer to caption Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption GAP-Net DGSMP DiffSCI DAUHST Ours (3stg) Ours (Mask)
@@ 4.4 Ablation Study
Table 4: Ablation study of CS-Mamba (3stg)’s components
@@
Baseline w/o SSM 0.37 37.41 0.964
with LE-SSM 0.77 38.83 0.975
with CS-SSM 0.57 37.88 0.968
with LE + CS SSM 0.97 39.10 0.977
Break-down Ablation. We conduct an ablation study by removing the LE-SSM block and CS-SSM block from the CS-Mamba network. The experiment results are shown in Table 4. Our goal is to explore the contributions of each mechanism to the results. The baseline barely adopts the parameter estimation, U-shaped framework, and GDFFN block without any SSM modules. Since [14] has demonstrated the effectiveness of GDFFN block, we don’t conduct an ablation study for it.
When we respectively employ LE-SSM and CS-SSM blocks, our model achieves 1.42 dB and 0.47 dB improvements in PSNR over the baseline, respectively. Employing both mechanisms simultaneously can further enhance the model’s performance. Here we need to point out that, in consideration of maintaining lightweight, we only utilized a single-direction scan in CS-SSM. In fact, we could also deploy a four-direction scanning similar to LE-SSM. However, each scan direction implies additional computational costs so we decided not to adopt this way.
Table 5: Comparisons of CS mechanism and common scanning.
@@
w/o Spectral SSM 38.83 0.975
with CS Mechanism 39.10 0.977
with Common Scanning 38.95 0.975
Cross-Scan Mechanism. To highlight the advantages of our cross-scan mechanism over common spectral scanning, we compared model performance under both approaches in our CS-Mamba (3stg) network. As shown in Table 5, common scanning provided a 0.12 dB improvement, whereas the proposed across-scanning mechanism yielded a significant 0.27 dB increase. Both methods were tested under single-scan conditions, with our method excelling due to its enhanced focus on adjacent bands and pixels, effectively capturing local features across spatial and spectral dimensions. These results underscore the effectiveness of our across-scanning approach.
@@
w/o Spatial SSM 37.88 0.968
with LE Machenism 38.83 0.975
with Global Scanning 38.66 0.974
Local-Enhancement Mechanism. We also conducted an ablation study of our LE-SSM mechanism in our CS-Mamba (3stg) to assess its impact. As shown in Table 6, starting from a baseline model with a PSNR of 37.88 dB (without spatial SSM), incorporating global four-direction scanning improved PSNR by 0.95 dB and SSIM by 0.007, demonstrating effective global spatial context modeling. Applying our LE mechanism along the spatial dimension further improved PSNR by 0.17 dB compared to global scanning, without introducing additional parameters or computational costs, unlike methods using CNN layers or spatial attention [24]. This enhancement highlights the significance of capturing local features in HSIs and validates the effectiveness of our LE mechanism.
@@ 5 Conclusion
In this paper, we propose a novel cross-scan SSM mechanism that more fully leverages the spectral and spatial local features of HSI images, achieving improved performance compared to the common SSM mechanism. Additionally, we identify and elaborate on noise issues encountered in real-world scene reconstruction, presenting a masked training method designed to address these challenges. This approach significantly mitigates the degradation problems in real-scene reconstruction that previous methods failed to resolve, enhancing the robustness of the model. Experiments in the main text and supplementary material validate the effectiveness of our method.